3/14/2018 3 Award: 10.00 points On June 30, 2018, Georgia-Atlantic, Inc., leased warehouse equipment from IC Leasing Corporation. a four-year lease term, payable each June 30 and December 31, with the first payment at June 30, 20 The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $604,152 over payment amounts. IC purchased the warehouse from Builders, Inc.. at a cost of $4.1 million. (FV of tu. pVof $1. EVA of $1, PVA of $1. EVAD of $1 and PVADOSI) (Use appropriate factor(s) from 18 Georgia-Atlantics incremental borrowing rate is 10% the same rate IC used to calculate lease the tables provided.) Required 1. Wh at pretax amounts related to the lease would IC report in its balance sheet at December 31 2018? 2. What pretax amounts related to the lease would IC report in its income statement for the year ended December 31, 2018? requirements, enter your answers in whole dollars and not in millions. Round your fin answers to ne 1. Pretax amount of net receivable 2. Pretax amount of interest revenue rev 01 08 2018 QC_ CS-113237 Hints Hint#1
Answer #1
1 Net recievable
Beginning balance on June 30. 4100000
Less:First payment on June 30.
(all are towards principal) 604152
Second payment on Dec 31.
(Note:1) 429360 1033512
Ending balance on Dec 31. 3066488
Notes:
1. interest payable=(4100000-604152)*10%*6/12=174792
Principal component=604152-174792=429360
2 Interest revenue
June 30. 0
Dec 31. (Note:1) 174792
Total 174792